OpenStudy (anonymous):

what interval is this increasing on -3sin^3 x (this is the derivative form)

OpenStudy (shayaan_mustafa):

HI jinnie :) do some magic :D lol so it is \[\large f'(x)=-3sin^3x\] Right friend?

OpenStudy (anonymous):

yes

OpenStudy (shayaan_mustafa):

OK remember rules now.. if \[\large f'(x)<0\] then function is decreasing. if\[\large f'(x)>0\]then function is increasing.

OpenStudy (anonymous):

yea i tried those already

OpenStudy (anonymous):

i have tried all these questions im posting. i cant get a result hence the reason im posting them

OpenStudy (anonymous):

if you know how to do this, can you just help me out

OpenStudy (anonymous):

its a hw question thats due soon and ive tried all i know

OpenStudy (anonymous):

@Jinnie is \(-3\sin^3(x)\) the funtion or the derivative of the function?

OpenStudy (anonymous):

sine is positive on \((0,\pi)\) and negative on \((\pi,2\pi)\)

OpenStudy (anonymous):

therefore \(-3\sin^3(x)\) is negative on \((0,\pi)\) and postive on \((\pi,2\pi)\)

OpenStudy (anonymous):

therefore your origninal function, whatever it is, is increasing on \((\pi,2\pi)\) and decreasing on \((0,\pi)\)

OpenStudy (anonymous):

thank you very much satellite. this is the derivative of -3sin^3(x) -9sin^2(x) cos(x) what would you say this is conave up and down on and the point of inflection. lastly for -3sin^3(x)...i got the local min to be pi, cant figure out the local max